OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (7, 2pi/3)

OpenStudy (anonymous):

Awesome... Polar coordinates \[\huge (r , \theta)\] Then your rectangular coordinates are given by \[\huge (x,y) = (r \cos \ \theta, r \sin \ \theta)\]

OpenStudy (anonymous):

@PeterPan so i would put \[(7\cos \theta) , (\frac{ 2\Pi }{ 3 }\sin \theta)\] into my calculator?

Parth (parthkohli):

Here, \(\theta = \dfrac{2\pi}{3}\). And don't do the commas.

OpenStudy (anonymous):

Nope... Remember, \[\huge (r, \theta) = (7, \frac{2\pi}{3})\]

OpenStudy (anonymous):

So your r is only ever 7 And your theta is only ever two-thirds of pi :)

Parth (parthkohli):

Figure out what \(7\cos \left(\frac{2\pi}3{}\right)\) is, then figure out what \(7\sin\left(\frac{2\pi}{3}\right)\) is. Don't forget the comma and parentheses.

Parth (parthkohli):

I didn't know Peter Pan was so good at mathematics. :-)

OpenStudy (anonymous):

When you have all the time in the world, sometimes, it pays to learn stuff :D

OpenStudy (anonymous):

@PeterPan and @ParthKohli oh wow, that makes so much more sense!! so then \[x=\frac{ -7 }{2} y=\frac{ 7\sqrt{3} }{2}\]

OpenStudy (anonymous):

Awesome work :)

OpenStudy (anonymous):

wow thank you both so much!! you both made it much easier than i was!! @PeterPan @ParthKohli

OpenStudy (anonymous):

Polar to Rectangular is such... Now the other way around, not as simple :P

OpenStudy (anonymous):

@PeterPan :( that's what i have to do now! can you help me with that also?

OpenStudy (anonymous):

I'm at your call :)

OpenStudy (anonymous):

@peterpan Find all polar coordinates of point P where P = \[( 1, \frac{ \Pi }{ 3 })\]

OpenStudy (anonymous):

Are you sure it's polar coordinates we want and not rectangular? The presence of a 'pi' in your coordinates is worth considering...

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

I think P is already in polar, and we're just trying to find which other polar coordinates are equivalent to it.

OpenStudy (anonymous):

@PeterPan i don't know how to do that.. this section that i'm doing is the toughest for me..

OpenStudy (anonymous):

Well, think of it this way... your theta, or your angle, if you rotate it one full rotation, or 2pi, it's basically the same angle, right?

OpenStudy (anonymous):

@PeterPan yes

OpenStudy (anonymous):

So you can add 2pi to it, and it stays the same, right?

OpenStudy (anonymous):

@PeterPan yes.. ?

OpenStudy (anonymous):

You can add it as many times as you want, right?

OpenStudy (anonymous):

@PeterPan yes

OpenStudy (anonymous):

So you can add 2npi, where n is any integer. Got that so far? :)

OpenStudy (anonymous):

@PeterPan i was just about to ask what n was, okay i understand that now!

OpenStudy (anonymous):

However, that's not the only way to rewrite your polar coordinates. You can also negate the radius (make it negative) and then, rotate the angle by a half-turn, or just pi (instead of 2pi)

OpenStudy (anonymous):

@PeterPan i understand that too

OpenStudy (anonymous):

and once you've added pi to it, you can once again add 2pi as many times as you want. So it's 2npi + pi, right?

OpenStudy (anonymous):

@PeterPan yes

OpenStudy (anonymous):

so it's (2n+1)pi, if you factor out the pi :P

OpenStudy (anonymous):

@PeterPan and that right there will allow me to find any coordinate on the plane?

OpenStudy (anonymous):

Well, to summarise, given our polar coordinates \[\large (r, \theta)\]It is always the same when we add any multiple of 2pi to the angle: \[\huge (r, \ \theta + 2n \pi)\] It is also the same if we negate the radius, and then rotate by a half turn by adding pi... \[\large (-r, \ \theta + \pi)\] But much like the first case, we can also keep on adding 2pi to this new angle and have it stay the same... \[\large (-r , \ \theta + \pi + 2n \pi)\] factoring out pi \[\huge (-r , \ \theta +(2n+1) \pi)\]

OpenStudy (anonymous):

@PeterPan omg okay, that makes so much sense! thank youu!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@PeterPan do you mind helping me with one more? :)

OpenStudy (anonymous):

Sure... let me have it :)

OpenStudy (anonymous):

@PeterPan Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

Okay, here's the thing about rectangular to polar... It's simple getting r You know how?

OpenStudy (anonymous):

@PeterPan so is this the opposite of what we doing earlier?

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

@PeterPan you would use \[r=\sqrt{x ^{2}+y ^{2}} , \theta=\tan^{-1} (\frac{ y }{ x })\]

OpenStudy (anonymous):

Yeah, that's good. Seems you already know what you're doing :) go ahead and plug in these formulas with x = 3 y = -3

OpenStudy (anonymous):

Not yet done, though, you're asked for two pairs of coordinates. Work out one pair, first, though

OpenStudy (anonymous):

@PeterPan well r would equal \[\sqrt{18}\] and \[\theta= -.785\] that's what i got..

OpenStudy (anonymous):

Remember, we're working with degrees this time. Try to redo the theta part.

OpenStudy (anonymous):

@PeterPan i would plug \[\tan^{-1}( \frac{ -3 }{ 3 })\] into my calculator correct?

OpenStudy (anonymous):

Yeah, but make sure it's set to degrees, and not radians.

OpenStudy (anonymous):

@peterpan i got -45 i forgot to change it into degrees! thanks

OpenStudy (anonymous):

Okay, so you got -45, but that doesn't really fall within the 0° ≤ θ < 360° restriction you wanted, so what to do?

OpenStudy (anonymous):

@PeterPan change it to positive! or make r negative

OpenStudy (anonymous):

change it to positive... how? By the way, there's no need to keep tagging me, it's like you're calling out my name every time you answer me :D

OpenStudy (anonymous):

oh, i wasn't sure if you would still get the notification! but okay but i'm honestly not too sure what to do since -45 isn't in the range...

OpenStudy (anonymous):

Well, remember, when we have θ, we can keep adding 2pi to it, and it basically stays the same? Well, when we have degrees, we can keep adding 360 to it, without it changing ! :)

OpenStudy (anonymous):

so we just add 2pi and it'll be postive right?

OpenStudy (anonymous):

This is in degrees, so you'll want to add 360 degrees...

OpenStudy (anonymous):

oh so it will be 315

OpenStudy (anonymous):

That's right :)

OpenStudy (anonymous):

But that's not the only pair of coordinates...

OpenStudy (anonymous):

there's another one..

OpenStudy (anonymous):

mhmm but before that, what was your first pair?

OpenStudy (anonymous):

\[(3\sqrt{2},315)\]

OpenStudy (anonymous):

lol okay, but remember what we said? It would be the same if we negate the radius (r) and then add (or subtract, it don't matter, really) pi to the angle. So, instead of pi, we add (or subtract) 180 degrees. Go ahead now :)

OpenStudy (anonymous):

it'll equal 135

OpenStudy (anonymous):

will there be a negative in front of \[(3\sqrt{2})\]

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

So what are your two pairs?

OpenStudy (anonymous):

(3sprt2, 315) and (-3sqrt2, 135) :)

OpenStudy (anonymous):

thank you so much!! i appreciate your help so much!!!

OpenStudy (anonymous):

No problem :)